OpenStudy (anonymous):

Determine if (-3/10 -2) is a solution to 10x-9y=15 . Please show all of your work.

OpenStudy (karatechopper):

10x-9y=15 10(-3/10)-9(-2)=15 -3+18=15 15=15 check!